OpenStudy (jaylelile):

I need help, please! Problem below:

OpenStudy (jaylelile):

\[\frac{ x+1 }{ 5 }-\frac{ x-6 }{ 3 }=\frac{ 1 }{ 3 }\] solve for x

OpenStudy (freckles):

you could clear the fractions by multiplying both sides by 5*3 aka 15

OpenStudy (freckles):

that should make the problem easier for you

OpenStudy (freckles):

let me see what you after doing this step please

OpenStudy (jaylelile):

I got x+1-x-6=1 then I made it 1x+1+6=1 The answer I got is x=-6

OpenStudy (mathmale):

Please check your answer by substituting it back into the original equation. Is that equation now true?

OpenStudy (mathmale):

Multiplying the original equation by the LCD (which is 15) results in 3(x+1) - 5)x-6) = 5, which differs from your result. Try solving this for x and checking your result. This will help you zero in on the correct approach.

OpenStudy (pawanyadav):

As LCM is multiply both side by 15 You get 15(x+1)/5 -15(x-6)/3=15/3 3(x+1)-5(x-6)=5 Now solve it by distribution..

OpenStudy (pawanyadav):

By distribution 3*x +3*1 - 5*x -5*(-6)=5 3x+3-5x+30=5 -2x+33=5 Can you solve this for x?

OpenStudy (jaylelile):

x=14 right?

OpenStudy (pawanyadav):

Yes right